Define stagnation temperature. Air leaves the compressor of a jet engine at a temperature of

150°C, a pressure of 5 bar and a velocity of 120m/s. Determine the isentropic stagnation

temperature and pressure. 


1
Expert's answer
2021-03-14T19:20:54-0400

"T=T_0+\frac{v^2}{c},"

"T=150+\frac{120^2}{1005}=164°C."


"p=p_0\cdot \frac{T}{T_0},"

"p=5\cdot \frac{164}{150}=5.5~bar."